# Structured Sparse Transition Matrices to Enable State Tracking in State-Space Models

### v2

@@ 3 PD Parametrization for Efficient and Expressive Transition Matrices
Prior research indicates that the expressiveness of time-varying SSMs can be significantly enhanced by relaxing the structural constraints of the transition matrix $A(u_{t})$ to allow for non-diagonal forms
[Merrill et al. 2024, Cirone et al. 2024, Terzić et al. 2025]. As noted by Terzić et al. [2025], a relaxation to arbitrary matrices prohibitively increases the compute cost of parallel scans from $\Theta(LN)$ to $\Theta(LN^{3})$. To circumvent these computational limitations, we instead propose to parameterize the transition matrices as the product $A(u_{t})=P(u_{t})D(u_{t})$, where $P(u_{t})$ is an input-dependent binary matrix in which each column has a single non-zero element, and $D(u_{t})$ is a complex-valued diagonal matrix.
@@
$\displaystyle s(u_{i})$ $\displaystyle=\text{softmax}(Su_{i})\in\Delta^{K-1}$
$\displaystyle M(u_{i})$ $\displaystyle=\sum_{k=1}^{K}s_{k}(u_{i})M_{k}\in\mathbb{R}^{N\times N}$
$\displaystyle P_{:,j}(u_{i})$ $\displaystyle=\text{hardmax}(M_{:,j}(u_{i}))\in\{0,1\}^{N}\text{ where }\text{hardmax(x)}_{i}:=\delta_{i,\arg\max_{j}x_{j}}$
The parametrization is motivated by Terzić et al. [2025], which has shown that models utilizing normalized Fan et al. [2024] convex combination of dense transition matrices can achieve perfect length generalization on FSA emulation tasks and are parameter-efficient compared to alternative proposals such as e.g. [Hasani et al. 2023, Fan et al. 2024, Merrill et al. 2024].
The $P$ and $D$ factors of our parametrization have complementary strengths for encoding FSAs into time-varying SSMs. The $P$ matrix enables emulating arbitrary FSAs, but the required dimension scales linearly with the number of states. For cyclic FSAs, which form a central building block of all solvable automata [Krohn and Rhodes 1965, Sarrof et al. 2024], complex diagonal matrices provide a more compact encoding compared to column one-hot matrices. Visual intuition is provided in Figure 2. Additionally, as proven in the following section, the diagonal matrices provide a guarantee for the system’s BIBO stability as the magnitude of each entry lies in $(0,1)$.
@@ 3.1 Approximate Gradients
This is reminiscent of the slope-annealed straight-through estimator [Paulus et al. 2021], yet we use stochasticity neither in the forward nor in the backward pass (ablations with stochastic categorical sampling in $P$ are reported in Appendix E). Note that during the forward pass we do not relax the hardmax, since doing so would break the sparsity essential for an efficient parallel scan. With tempered softmax in the low-temperature limit, the above expression becomes exact.
@@ 3.2 Algebraic Structure of the PD Parametrization
In this section, we formalize the set of transition matrices $\mathbb{H}^{N\times N}$ that are decomposable into a binary one-hot column matrix $P$ and a complex diagonal matrix $D$. Note that our $PD$ matrices exhibit such a column one-hot structure. Hence, all of the subsequent statements made for the abstract class $\mathbb{H}$ immediately also hold for our matrix parametrization. All proofs are in Appendix C.
Definition 1 (Monoid of One-hot Column Matrices).
Let $\mathbb{H}^{M\times N}:=\{A\in\mathbb{C}^{M\times N}:\forall i\ \|A_{:,i}\|_{0}=1\}$ where $\|x\|_{0}$ denotes the $\ell^{0}$-"norm" counting the number of non-zero entries in $x$. $H^{N\times N}$ forms a monoid under matrix multiplication.
@@
Closure under multiplication is essential for efficient chained matrix multiplication via parallel scans, because matrix multiplication in $\mathbb{H}^{N\times N}$ can be implemented in $\Theta(N)$ instead of the usual $\Theta(N^{3})$, as visualized in Figure 3.
@@ 3.3 Stability and Expressivity of the PD Parametrization
ov Let $\varepsilon\in(0,1]$ and consider the state transition $x_{t}=A_{t}x_{t-1}+b_{t}$ with $A_{t}\in\mathbb{H}^{N\times N}:\|A_{t}\|_{\infty}\leq 1-\varepsilon$. Let further $\|x_{0}\|_{2}\leq B$ and $\|b_{t}\|_{2}\leq B$ for $B\in\mathbb{R}_{+}$. Then it holds that
@@
As an immediate consequence of the mapping from FSAs to SSMs described in Section 2.2, Proposition 2 holds:
@@
Not only can PD-SSMs represent any FSAs, they do so with the (almost) smallest state size possible, i.e., PD-SSMs are maximally expressive for regular languages.
@@ 4.1 Runtime Measurements
We start by benchmarking the effective runtime for performing a forward pass of a single SSM layer with variable state size on an NVIDIA A100-80GB GPU. The diagonal SSM is our model without the $P$ matrices, whereas SD-SSM uses unstructured transition matrices [Terzić et al. 2025]. As shown in Figure 4, PD-SSM scales significantly better than SD-SSM. PD-SSM’s higher runtime compared to diagonal SSM stems primarily from the additional operations in the generation of P: the simultaneous soft selection of transition matrices $M_{i}$ across the sequence during parallel execution incurs additional compute and memory. Table S4 in Appendix E confirms these findings. [Uncaptioned image] Figure 4: Runtime comparison of single-layer SSMs with different transition matrix structures as a function of state size with sequence length $L=1024$.
@@ 4.2 FSA Emulation
We first evaluate our model on a set of automaton state tracking tasks, originally introduced in Delétang et al. [2023]. The four tasks correspond to four finite-state automata of various complexities. The benchmark is centered at evaluating the length generalization of the models, which serves as a proxy of the model having learned the correct algorithmic abstraction, avoiding fixed-length shortcut solutions Liu et al. [2023]. Concretely, the models are trained for 100,000 steps on a dataset of randomly sampled sequences of symbolic inputs of length 3 to 40, and are evaluated on sequences of length 40–256. We extend the set of results from Walker et al. [2025], which evaluates each model under only one varying hyperparameter choice, state dimensionality of 128 or 512, reporting the best obtained result. We instead fix the dimensionality to 128, finding that it is sufficient for high performance. The other hyperparameters are fixed for each model. All models are evaluated using two layers. Our Table 2 reports the mean and standard deviation of the best validation accuracy of five randomly initialized models. The set of evaluated models consists of recurrent and parallelizable models, where we bold and underline the best and second-best parallelizable model, respectively. Excluding the Transformer, the parallelizable models can be interpreted as SSMs with various transition matrix structures, namely diagonal (Gu and Dao [2023], Walker et al. [2025] and our $\mathbb{C}$ diagonal model defined by setting $P(u_{t})$ to identity), diagonal plus low-rank (Schlag et al. [2021a], Yang et al. [2024], Grazzi et al. [2024]), product of diagonal plus low-rank [Walker et al. 2025, Peng et al. 2025, Siems et al. 2025], as well as alternative variants including block-diagonal, a Walsh-Hadamard matrix modulated by an input-dependent diagonal matrix, and a mixture of diagonal and dense matrices Walker et al. [2025]. As we can read from the table, our model performs best, with a significant margin over the second-best method which utilizes dense matrices every $16^{\text{th}}$ transition.
@@
Sequential
LSTM 100.0 ± 0.0 100.0 ± 0.0 99.9 ± 0.1 100.0 ± 0.0 100.0
sLSTM 39.2 ± 4.3 100.0 ± 0.0 34.8 ± 1.5 100.0 ± 0.0 68.5
xLSTM[1:1] 56.6 ± 3.6 100.0 ± 0.0 34.4 ± 1.4 100.0 ± 0.0 72.8
@@
Transformer 24.4 ± 0.5 90.4 ± 10.4 23.6 ± 0.7 52.2 ± 0.4 47.7
Mamba 50.7 ± 7.8 100.0 ± 0.0 38.1 ± 6.0 54.5 ± 1.5 60.8
D–LNCDE 73.5 ± 4.7 100.0 ± 0.0 20.8 ± 0.3 91.9 ± 14.5 71.6
$\mathbf{\mathbb{C}}$ Diag. 90.4 ± 4.3 82.4 ± 4.8 59.9 ± 27.0 61.8 ± 6.8 73.6
DeltaNet 49.0 ± 5.4 100.0 ± 0.0 44.2 ± 8.0 58.4 ± 1.1 62.9
Gated DeltaNet 55.9 ± 7.8 100.0 ± 0.0 49.9 ± 7.5 56.6 ± 0.7 65.6
DeltaNet[-1, 1] 47.3 ± 6.0 100.0 ± 0.0 63.9 ± 7.2 97.2 ± 1.7 77.1
RWKV-7 37.8 ± 5.0 88.1 ± 14.2 41.6 ± 13.8 51.0 ± 0.3 54.6
DPLR4–LNCDE 83.9 ± 20.1 97.7 ± 5.2 26.0 ± 0.7 79.6 ± 20.3 71.8
Gated DeltaProduct[-1,1] 61.4 ± 12.0 99.6 ± 0.6 58.0 ± 20.0 98.8 ± 1.5 79.5
WH–LNCDE 66.7 ± 7.9 87.5 ± 17.2 23.9 ± 1.3 62.1 ± 2.2 60.1
BD4–LNCDE 100.0 ± 0.0 91.3 ± 12.2 45.6 ± 7.8 88.4 ± 13.9 81.3
D–DE16–LNCDE 77.3 ± 21.3 81.4 ± 9.9 94.5 ± 5.2 81.0 ± 18.1 83.6
PD-SSM 99.5 ± 0.7 99.7 ± 0.3 96.2 ± 3.4 99.9 ± 0.1 98.8
@@
Table 2: Average and standard deviation of validation accuracy across 5 seeds for a range of models on FSA emulation tasks. The baseline results are taken from [Walker et al. 2025].
We further analyse the method on two non-solvable groups, $A_{5}$ and $S_{5}$. For both groups, consisting of 60 and 120 permutations respectively, all of the group elements can be generated using only two permutations, corresponding to two transition matrices as per the mapping in Secion 2.2. To increase the connectivity of the resulting automaton’s states, we introduce additional randomly selected permutations. We compare a PD-SSM with $K=32$ against two layers of the complex diagonal model with $A(u_{t})=D(u_{t})$, and one or two layers of a DPLR model, Gated DeltaProduct [-1,1] with $n_{h}=4$ [Siems et al. 2025], with state dimension $128$. We perform a learning rate grid search and train the models for 100,000 steps with batch size 256. We train the models on sequences of length up to 40 and report the best validation accuracy on sequences of length up to 256 in Table 3.
@@
Table 3: Best validation accuracy on longer sequences across 3 random seeds. Each task is defined in terms of a non-solvable algebraic group ($A_{5}$ or $S_{5}$) generated with redundant set of permutations, the cardinality of which is the number $n$ in each tuple $(A_{5}\text{ or }S_{5},n)$.
The complex diagonal model expectedly fails on the simplest example $(A_{5},2)$. Gated DeltaNet with $n_{h}=4$ exhibits degrading accuracy with higher number of permutations, which cannot be recovered by stacking two layers of the model. PD-SSM maintains full accuracy for all automata except $S_{5}$ with $32$ generating permutations.
@@ 4.3 Multivariate Time-Series Classification
We next provide an evaluation on multivariate time-series classification. We evaluate our model on a subset of the University of East Anglia (UEA) Multivariate Time-Series Classification Archive (UEA-MTSCA) [Dau et al. 2019], extending on the results from Walker et al. [2024], Rusch and Rus [2025]. We consider six tasks from the archive previously selected due to their long sequence lengths, which range from 800 to over 17,000. Conforming to the evaluation methodology and data splits defined in Rusch and Rus [2025], Walker et al. [2024], we report the average and standard deviation of the test accuracy over five random initializations, with the hyperparameter grid conforming to that of [Rusch and Rus 2025] with the exception of the state size, which in our case is selected from $\{16,64,128\}$ as opposed to $\{16,64,256\}$ of the baselines. The set of evaluated models includes neural controlled differential equations [Walker et al. 2024], LTI SSMs [Smith et al. 2023, Orvieto et al. 2023], time-varying SSMs [Gu and Dao 2023], as well as an SSM paradigm based on a system of forced harmonic oscillators [Rusch and Rus 2025]. The results are reported in Table 4. As we can see, our model maintains very high accuracy on this set of tasks, achieving the second-best average.
@@
Table 4: Mean and standard deviation of test accuracies across 5 seeds on selected long-sequence UEA time-series classification datasets as per [Rusch and Rus 2025].

### v3

@@ 3 PD Parametrization for Efficient and Expressive Transition Matrices
Prior research indicates that the expressiveness of time-varying SSMs can be significantly enhanced by relaxing the structural constraints of the transition matrix $A(u_{t})$ to allow for non-diagonal structures. [merrill_illusion_2024, cirone_theoretical_2024, terzic2025sdssm]. As noted by terzic2025sdssm, a relaxation to arbitrary matrices prohibitively increases the compute cost of parallel scans from $\Theta(LN)$ to $\Theta(LN^{3})$. To circumvent these limitations, we propose to parameterize the transition matrices as the product $A(u_{t})=P(u_{t})D(u_{t})$, where $P(u_{t})$ is a binary matrix in which each column has a single non-zero element, and $D(u_{t})$ is a complex-valued diagonal matrix.
@@
$\displaystyle s(u_{t})$ $\displaystyle=\text{softmax}(Su_{t})\in\Delta^{K-1}$
$\displaystyle M(u_{t})$ $\displaystyle=\sum_{k=1}^{K}s_{k}(u_{t})M_{k}\in\mathbb{R}^{N\times N}$
$\displaystyle P_{:,j}(u_{t})$ $\displaystyle=\text{hardmax}(M_{:,j}(u_{t}))\in\{0,1\}^{N}\text{ where }\text{hardmax(x)}_{i}:=\delta_{i,\arg\max_{j}x_{j}}$
The parametrization is motivated by terzic2025sdssm, which has shown that models utilizing normalized fan_advancing_2024 convex combinations of dense transition matrices can achieve perfect length generalization on FSA emulation tasks and are parameter-efficient compared to alternative proposals such as e.g. [hasani2023liquid, fan_advancing_2024, merrill_illusion_2024].
The $P$ and $D$ factors of our parametrization have complementary strengths for encoding automata into time-varying SSMs. The $P$ matrix enables emulating any FSA, but the required dimension scales linearly with the number of states. For cyclic automata, which form a central building block of all solvable automata [krohn_algebraic_1965, sarrof2024expressivecapacitystatespace], complex diagonal matrices provide a more compact encoding compared to column one-hot matrices. Visual intuition is provided in Figure 2. Additionally, as proven in the following section, the diagonal matrices provide a guarantee for the system’s BIBO stability as the magnitude of each entry lies in $(0,1)$.
@@ 3.1 Surrogate Gradients
This is reminiscent of the slope-annealed straight-through estimator [bengio_stochastic_2013, paulus_rao_2021], yet we use stochasticity neither in the forward nor in the backward pass (ablations with stochastic categorical sampling in $P$ are reported in Appendix E). Note that during the forward pass we do not relax the hardmax, since doing so would break the sparsity essential for an efficient parallel scan. With tempered softmax in the low-temperature limit, the above expression becomes exact.
@@ 3.2 Algebraic Structure of the PD Parametrization
In this section, we formalize the set of transition matrices $\mathbb{H}^{N\times N}$ that are decomposable into a binary one-hot column matrix $P$ and a complex diagonal matrix $D$. Note that our $PD$ matrices exhibit such a column one-hot structure. Hence, all of the subsequent statements made for the set $\mathbb{H}$ immediately also hold for our matrix parametrization. All proofs are in Appendix C.
Definition 1 (One-hot Column Matrices).
Let $\mathbb{H}^{M\times N}:=\{A\in\mathbb{C}^{M\times N}:\forall i\ \|A_{:,i}\|_{0}=1\}$ where $\|x\|_{0}$ denotes the $\ell^{0}$-"norm" counting the number of non-zero entries in $x$.
@@
Closure under multiplication is essential for efficient chained matrix multiplication via parallel scans, because matrix multiplication in $\mathbb{H}^{N\times N}$ can be implemented in $\Theta(N)$ instead of the usual $\Theta(N^{3})$, as formalized by Property 2 and visualized in Figure 3.
@@ 3.3 Stability and Expressivity of the PD Parametrization
Let $\varepsilon\in(0,1]$ and consider the state transition $x_{t}=A_{t}x_{t-1}+b_{t}$ with $A_{t}\in\mathbb{H}^{N\times N}:\|A_{t}\|_{\infty}\leq 1-\varepsilon$. Let further $\|x_{0}\|_{2}\leq B$ and $\|b_{t}\|_{2}\leq B$ for $B\in\mathbb{R}_{+}$. Then it holds that
@@
As a direct consequence of the FSA to SSM mapping described in Section 2.2, Proposition 2 holds:
@@
Not only can PD-SSMs represent any FSA, they do so with the (almost) smallest state size possible, i.e., PD-SSMs are maximally expressive for regular languages.
@@ 4.1 Runtime Measurements
We first measure how the runtime of a single-layer SSM scales as a function of the transition matrix structure as well as the hidden dimension on an NVIDIA A100-80GB GPU. We compare PD-SSM with the unstructured (dense) real-valued SD-SSM [terzic2025sdssm], as well as a diagonal variant of PD-SSM that sets $A(u_{t})=D(u_{t})\in\mathbb{C}^{N_{c}}$. To equalize the effective state size, the state dimensionality $N_{r}$ of the real-valued dense SSM is twice that of the complex-valued models $N_{c}$. The embedding size is scaled as $D=N_{r}=2N_{c}$. In Figure 4 we see that PD-SSM scales significantly better than the dense SSM, with a $71\times$ speed-up at $D=5632$. At this dimension, the diagonal SSM is $7\times$ faster than PD-SSM. The PD model’s higher runtime mainly stems from the additional operations in the generation of P matrices. For more details and results under different settings, see Appendix E. [Uncaptioned image] Figure 4: Runtimes of single-layer SSMs with varying dimension and sequence length $64$.
@@ 4.2 FSA Emulation
We first evaluate our model on a set of automaton state tracking tasks, originally introduced in deletang_neural_2023. The four tasks correspond to four finite-state automata of various complexities. The benchmark is centered at evaluating the length generalization of the models, which serves as a proxy of the model having learned the correct algorithmic abstraction, avoiding fixed-length shortcut solutions liu_transformers_2023. Concretely, the models are trained for 100,000 steps on randomly sampled sequences of inputs of length 3 to 40, and are evaluated on sequences of length 40–256. We extend the set of results from walker_2025_structured which evaluates each model under a single varying hyperparameter choice, state dimensionality of 128 or 512. We instead fix the dimensionality to 128, finding that it is sufficient for high performance. Table 2 reports the mean and standard deviation of the best validation accuracy of five randomly initialized models. The set of evaluated models consists of recurrent and parallelizable models, where we bold and underline the best and second-best parallelizable model, respectively. Excluding the Transformer, the parallelizable models can be interpreted as SSMs with various transition matrix structures, namely diagonal (gu_mamba_2023, walker_2025_structured and our $\mathbb{C}$ diagonal model defined by setting $P(u_{t})$ to identity), diagonal plus low-rank (schlag2021linear, yang2024parallelizing, grazzi2024unlocking), product of diagonal plus low-rank [walker_2025_structured, peng_2025_rwkv, siems_2025_deltaproduct], as well as alternative variants including block-diagonal, a Walsh-Hadamard matrix modulated by an input-dependent diagonal matrix, and a mixture of diagonal and dense matrices walker_2025_structured. As we can read from the table, our model performs best, with a significant margin over the second-best method.
@@
Recurrent
LSTM 100.0 ± 0.0 100.0 ± 0.0 99.9 ± 0.1 100.0 ± 0.0 100.0 ± 0.0
sLSTM 32.5 ± 0.4 100.0 ± 0.0 27.7 ± 0.6 100.0 ± 0.0 65.1 ± 0.2
xLSTM[1:1] 53.5 ± 5.6 99.0 ± 1.9 29.3 ± 1.6 100.0 ± 0.0 70.5 ± 1.5
@@
Transformer 24.4 ± 0.5 90.4 ± 10.4 23.6 ± 0.7 52.2 ± 0.4 47.7 ± 2.6
Mamba 48.4 ± 2.2 100.0 ± 0.0 33.1 ± 6.6 54.2 ± 2.1 58.9 ± 1.8
D-SLiCE 69.5 ± 6.3 100.0 ± 0.0 20.9 ± 0.1 100.0 ± 0.0 72.6 ± 1.6
$\mathbf{\mathbb{C}}$ Diag. 90.4 ± 4.3 82.4 ± 4.8 59.9 ± 27.0 61.8 ± 6.8 73.6 ± 7.1
DeltaNet 49.8 ± 4.7 100.0 ± 0.0 42.2 ± 4.8 57.8 ± 0.8 62.5 ± 1.7
DeltaNet[-1,1] 46.7 ± 6.1 100.0 ± 0.0 66.4 ± 8.8 97.7 ± 2.0 77.7 ± 2.7
Gated DeltaNet 53.8 ± 8.8 100.0 ± 0.0 42.8 ± 8.2 56.5 ± 1.9 63.3 ± 3.0
Gated DeltaProduct[-1,1] 46.3 ± 6.6 100.0 ± 0.0 78.4 ± 10.9 98.0 ± 1.4 80.7 ± 3.2
RWKV-7 37.8 ± 5.0 88.1 ± 14.2 39.5 ± 6.1 51.1 ± 0.3 54.1 ± 4.1
[…]
WH-SLiCE 69.7 ± 8.8 93.1 ± 13.9 23.8 ± 1.1 71.4 ± 12.9 64.5 ± 5.2
[…]
PD-SSM 99.5 ± 0.7 99.7 ± 0.3 96.2 ± 3.4 99.9 ± 0.1 98.8 ± 0.9
@@
Table 2: Average and standard deviation of validation accuracy across 5 seeds for a range of models on FSA emulation tasks. The baseline results are taken from [walker_2025_structured].
We further analyse the method on two non-solvable groups, $A_{5}$ and $S_{5}$. For both groups, consisting of 60 and 120 permutations respectively, all of the group elements can be generated using only two permutations corresponding to two transition matrices as per the mapping in Secion 2.2. To increase the connectivity of the resulting automaton’s states, we introduce additional randomly selected permutations. We compare a PD-SSM with $K=32$ against two layers of the complex diagonal model with $A(u_{t})=D(u_{t})$, and one or two layers of a DPLR model, Gated DeltaProduct [-1,1] with $n_{h}=4$ [siems_2025_deltaproduct], with state dimension $128$. We perform a learning rate grid search and train the models for 100,000 steps with batch size 256. We train the models on sequences of length up to 40 and report the best validation accuracy on sequences of length up to 40-256 in Table 3.
@@
Table 3: Best validation accuracy (%) on longer sequences across 3 random seeds. Each task is defined in terms of a non-solvable algebraic group ($A_{5}$ or $S_{5}$) generated with redundant set of permutations, the cardinality of which is the number $n$ in each tuple $(A_{5}\text{ or }S_{5},n)$.
As predicted by the theory, the complex diagonal model fails on the simplest task, $(A_{5},2)$. Gated DeltaProduct with $n_{h}=4$ exhibits degrading accuracy with a higher number of permutations, which cannot be recovered by stacking two layers of the model. PD-SSM maintains full accuracy for all automata with a single layer, excepting $S_{5}$ with $32$ generating permutations.
@@ 4.3 Multivariate Time-Series Classification
We next provide an evaluation on multivariate time-series classification. We evaluate our model on a subset of the University of East Anglia (UEA) Multivariate Time-Series Classification Archive (UEA-MTSCA) [uea_time_series], extending on the results from walker_2024_logncde, rusch2025oscillatory. We consider six tasks from the archive previously selected due to their long sequence lengths, which range from around 400 to over 17,000. Conforming to the evaluation methodology and data splits defined in rusch2025oscillatory, walker_2024_logncde, we report the average and standard deviation of the test accuracy across five random initializations, with the hyperparameter grid conforming to that of [rusch2025oscillatory] with the exception of the state size, which in our case is selected from $\{16,64,128\}$ as opposed to $\{16,64,256\}$ of the baselines. The set of evaluated models includes neural controlled differential equations [walker_2024_logncde], LTI SSMs [smith_simplified_2023, orvieto_resurrecting_2023], time-varying SSMs [gu_mamba_2023], as well as an SSM paradigm based on a system of forced harmonic oscillators [rusch2025oscillatory]. The results are reported in Table 4. As we can see, our model maintains very high accuracy on this set of tasks, achieving an estimated mean accuracy that is within standard error of the state-of-the art SSM.
@@
Table 4: Mean and standard deviation of test accuracies across 5 seeds on selected long-sequence UEA time-series classification datasets as per [rusch2025oscillatory].